# Title keyword 5 for citation Title for citation Title for citation

### v2

@@ 4.1.1 Connection for AI
For emerging services like AI, the current network’s support in communication is yet to be clarified, necessitating enhanced design to ensure that future 6G networks can better meet the new demands of AI services from a connectivity perspective. Based on the types of AI service-related data carried by communication connections and the principle of control and user plane separation [161], the data can be naturally categorized into two major groups: AI signaling and AI data. Consequently, connection for AI embodies providing connectivity support for AI signaling and data.
@@
AI Data: 1) The AI data is used to transfer AI input data, generally referred to as input vectors for invoking an AI model or represented as inputs in an AI algorithm. 2) It is also used for transmitting AI outputs, generally called the output of the invoked AI model/algorithm. 3) It can be further used for transmitting AI models, e.g., transmitting a neural network model usually means transmitting its parameter weights (which can be represented as vectors) and its network structure (which can be represented as scalars).
For AI business scenarios and related connectivity performance requirements, 3GPP TS 22.261 (based on the conclusions of TR 22.874) also has some KPI requirement level analysis and results for 5G systems, mainly on the impact of traditional connectivity performance metrics, such as throughput, latency, reliability, etc., i.e., the current 3GPP standard only analyzes the impact of the 5G system (mainly the core network portion) in terms of how it supports different AI services/applications from the functional and performance perspectives. From an overall end-to-end process analysis perspective, AI signaling and AI data involved in AI services/applications necessitate transmission via the air interface. However, in the current standards and architectures, no relevant modules guide the transmission process or standardize the transmission content. As a result, AI services, as an emerging service type under the vision of intelligence embedded, inherently demand that the network provide better support for AI based on connection.
@@ 4.1.2 Data for AI
Future mobile communication networks will continuously generate massive amounts of data with diverse types, and data is precisely one of the core production factors for AI, providing a solid data foundation for 6G’s endogenous intelligent AI services [171]. Leveraging this vast array of multi-category, multi-modal data to support a broad spectrum of AI services realizes the vision of 6G’s “ubiquitous intelligence”. Consequently, data for AI aims to efficiently support end-to-end data collection, transmission, storage, and sharing, addressing how to facilitate, expedite, and securely provide data to AI functions within or external to the 6G network. This endeavor strives to reduce the difficulty in data acquisition, enhance data service efficiency, and facilitate the realization of 6G AI while adhering to privacy and security laws and regulations. Depending on the potential functional scope, the data support provided by data for AI should encompass five components:
@@
Data Collection/Distribution. This component provides a basic publish-subscribe mechanism for data producers and consumers, enhancing data collection/distribution efficiency and supplying samples for large-scale training required by AI models to achieve optimal results.
@@
Data Security and Privacy. This component leverages security and privacy protection technologies to provide high-quality, trusted data services tailored to user and network needs, ensuring user and network privacy protection and data security, immutability, and traceability. By safeguarding data privacy, it also protects the security of AI models during training and data transmission.
@@
Data Analysis. This component utilizes AI models, algorithms, and knowledge to provide statistical information, predictive insights, network anomaly analysis, and optimization suggestions, enhancing the data consumption experience for internal and external network functions and related intelligent applications.
@@
Data Preprocessing. This component applies generic tools for format conversion, noise reduction, feature extraction, and other preprocessing tasks to collected data, fulfilling AI model training requirements for input data and facilitating the management of multi-dimensional complex data.
@@
Data Storage. This component stores and retrieves collected data, providing storage support for data services related to data security and privacy, data analysis, or data preprocessing, and further establishing a data foundation to support AI services.
Through standardized data service functions, a complete data service process can be established within the 6G network, providing a robust foundation for AI at the data level.
@@ 4.1.3 Computation for AI
Computation is one of the core production factors of AI. In future network architectures, computational power will be ubiquitous across various infrastructures, including central clouds, edge clouds, network devices, and even end-user devices, providing a solid computational foundation for the diversification of AI services across all industries in 6G [172, 173]. This computational power, along with the AI algorithms or functionalities attached, serves the network or devices to improve performance and optimize network operations but is also potentially exploited through unified interfaces to serve upper-layer applications. Consequently, computation for AI focuses on how the widely available computational power in the network can support AI services. This computational power can be broadly categorized into three types: dedicated computing power for network elements, distributed external computing power, and endogenous computing power within the network.
Network Element Dedicated Computing Power (for AI): This type of computing power typically consists of computing and storage units composed of specialized processors and programmable devices. It serves network elements in mobile communication networks (BSs or core networks). It is primarily used for customized AI application services (i.e., AI4NET) to enhance communication performance or optimize network operations, like CSI feedback, channel estimation, and beamforming. Due to resource restrictions, it cannot support AI services and applications requiring large-scale computations and training. It is challenging to implement for third-party applications.
Distributed External Computing Power (for AI): This type of computing power typically exists in the form of distributed Edge Computing / Multi-access Edge Computing (MEC), primarily utilizing general-purpose CPUs and potentially including high-performance GPUs and programmable acceleration cards. External computing power shifts computations from centralized data centers to the edge of the access network, enabling network optimization and supporting industry applications with high computational demands and stringent latency requirements, such as video acceleration and AR/VR scenarios. Due to its external nature, it requires management through unified network functions, introducing some latency that may impact extremely latency-sensitive AI services and applications. Moreover, this external overlay approach may not optimize resource utilization, hindering efficient construction, deployment, and usage of AI services.
@@ 4.1.4 Security for AI
Some security vulnerabilities remained after the design of 2G-5G systems, and traditional security measures often failed to ensure safety. In recent years, trusted endogenous security has emerged as a new security approach for 6G, characterized by four main features: collaboration, intelligent proactive defense, trustworthiness, and privacy protection. The concept of trusted endogenous security also applies to ensuring the security of the AI plane.
@@
Trustworthiness: 6G will further evolve towards trustworthiness capabilities, achieving a comprehensive endogenous security system encompassing Security (i.e., confidentiality, integrity, availability, CIA), Reliability, Resilience, and Safety [175]. This can provide a trusted ecosystem for NET4AI. Furthermore, flexible and reliable access control[176] and effective identity authentication[177] can, to some extent, prevent unauthorized access from impacting the endogenous trusted environment of NET4AI.
Privacy Protection: The deep integration of AI, big data, and 6G networks will exacerbate the challenges of data privacy protection. NET4AI must handle massive business and personal data while preventing privacy leakage. Emerging distributed machine learning, while achieving distributed ubiquitous intelligence and partially avoiding privacy leakage caused by AI model training, still has security risks [178]. Therefore, NET4AI will integrate distributed machine learning with existing privacy protection methods at multiple points to enhance data privacy protection and build an efficient and secure data ecosystem.
@@ 4.1.5 Orchestration for AI
Fault Self-Healing and Intelligent Operations and Maintenance (O&\&&M): By integrating monitoring, alerting, and troubleshooting functions, the orchestrator promptly identifies and locates faults within the network, automatically triggering recovery mechanisms. This fault self-healing capability minimizes service disruptions for AI services, enhancing service availability and reliability. Additionally, the orchestration facilitates intelligent network O&M, empowering network administrators to manage network resources and AI services more efficiently, reducing O&M costs, and improving efficiency.
@@ 4.2 Key Enabling Technologies
This section describes the key enabling technologies for Net4AI. We compile the key enabling technologies for five planes in 6G networks and present them in Figure 17. In the following, we focus on enabling technologies closely related to the needs of AI, introducing their research background, research significance, and research status.
@@ 4.2.1 Distributed Intelligence and Federated Learning
The NET4AI architecture must support low-latency, high-reliability communication computing requirements and increasing privacy concerns. Distributed intelligence and federated learning (FL) are the key technologies that can solve this problem. Traditional centralized architectures in wireless networks can no longer meet low-latency, high-reliability communication and computation demands. Additionally, conventional centralized architectures are inefficient in supporting the ubiquitous intelligence required by future wireless networks. This trend towards deploying intelligent decentralized elements has become apparent, with intelligence gradually shifting from centralized to distributed systems. Therefore, introducing distributed intelligent computing architectures is necessary to fully utilize the multi-dimensional data and computing resources held by user terminals and nodes, enabling intelligent connectivity.
@@
Wireless distributed intelligence refers to organizing AI tasks in a distributed manner within wireless networks, incorporating collaborative AI and machine learning without uploading all raw data to a central cloud. This approach alleviates the network bandwidth pressure and reduces the central cloud’s computational burden. Moreover, as the computing power of smartphones and IoT devices increases and users become more concerned about data privacy, mobile devices can handle more AI computations and training. This proximity of AI to users and data allows for faster iteration and upgrade cycles, efficiently supporting the pervasive intrinsic intelligence of 6G networks.
@@
Distributed intelligence is a key technology for achieving the intrinsic intelligence of 6G networks. It adopts a decentralized architecture, dispersing network functions and resources across multiple nodes or devices, eliminating the dependence on a single central node to control the entire network. Nodes in a distributed network possess a certain degree of autonomy, enabling them to make decisions and execute tasks independently without frequent communication or coordination with other nodes. This autonomy enhances network flexibility, allowing it to better adapt to various environmental and demand changes. Data is typically stored across multiple nodes in a distributed network, and computational tasks can be distributed among these nodes for processing. This distributed data storage and processing method improves network scalability, accelerates the training and inference processes of large-scale models, and enhances the efficiency of AI systems. Distributed intelligence also supports online learning and continuous optimization by collecting and processing data in real-time, continuously updating models, and improving AI system performance and accuracy. By leveraging the potential of large-scale distributed edge devices, distributed AI technology can overcome the bottlenecks of centralized AI, reduce communication costs, and address data privacy concerns. Distributed AI and swarm intelligence will build a new intelligent ecosystem for 6G. As shown in Figure 18, the wireless distributed intelligence architectures widely recognized in academia and industry include :
Figure 18: FL, Multi-Agent Reinforcement Learning and Split Learning
Horizontal FL: Also known as sample-based FL, this can be applied in scenarios where the datasets of the various participants in FL have the same feature space but different sample spaces. Horizontal FL enhances privacy by training AI models on local devices and sharing only model parameters rather than raw data. Each node trains the model locally during training and periodically uploads model weights. The central node integrates these weights to form federated model weights, which are then sent back to the nodes for the next training cycle or local inference. FL ensures that user data remains local, transmitting only weights/gradients to achieve global training, thereby enhancing privacy.
@@
Through horizontal federation learning, different participants can jointly train a more powerful AI model based on local data sets while protecting privacy and improving the accuracy and generalization of the model. Additionally, horizontal FL enables real-time learning on distributed edge devices, allowing for timely model parameter updates to constantly changing environments and data. This is crucial for AI applications that require quick responsiveness and strong adaptability. It also supports collaboration across devices with varying computational capabilities and data distributions, catering to the diverse needs of 6G network terminals. Personalized model training on local devices enhances user experience by offering more individualized services.
The main application areas of horizontal FL include smart healthcare, personalized services, and intelligent manufacturing[180]. The authors of [181] provide the framework for an advanced blockchain-orchestrated machine learning system for privacy-preserving FL in medicine and a new utility in health. The authors of [182] combined blockchain and FL to accelerate AI model training in healthcare and improve model robustness. The authors of [183] introduced personalized FL (PFL) methods with theoretical performance derivations, modeling user personalization and data distribution. The authors of [184] categorized PFL techniques based on key challenges and strategies, highlighting critical ideas and future research directions in PFL architecture, realistic benchmarks, and trustworthy methods. The authors of [185] propose a new FL framework for cross-domain prediction for intelligent manufacturing, designed to quickly adapt to new products, processes, and applications with limited training data.
@@
Vertical FL: Vertical FL refers to a setup where datasets have different features but share the same samples/users. It is suitable for scenarios where there is a high overlap of users among participants but low or no overlap in feature space by extracting parts where participants have the same users but not the same features for joint training. In the training process, it is first necessary to sample align each participant’s data to obtain overlapping user data, then train the model locally and upload the parameters to the server [186]. The difference between vertical and horizontal FL is that horizontal FL emphasizes cooperation and sharing between different data sources. In contrast, vertical FL focuses on collaboration and sharing within varying levels of the same data source.
@@
The main application areas of vertical FL include recommendation systems, smart healthcare, and smart transportation. The authors of [187] propose a method based on clustering and latent factor models under the vertical federated recommendation system. Considering the diversity of a large number of different users in each participant, users are clustered to reduce the dimension of the matrix and improve the accuracy of user recommendations. The authors of [188] describe the idea of Intention-Hiding Vertical FL and illustrate a framework to achieve this privacy-preserving goal. The authors of [189] propose a Federated Deep Learning based on the Spatial-Temporal Long and Short-Term Networks algorithm to predict traffic flow by utilizing observed historical traffic data. An additive homomorphic encryption method based on vertical FL trains local TFP models and shares short-term spatio-temporal hiding information.

### v3

@@ 3.1.1 Connection for AI
For emerging services like AI, the current network’s support in communication is yet to be clarified, necessitating enhanced design to ensure that future 6G networks can better meet the new demands of AI services from a connectivity perspective. Considering the control plane and user plane separation principle [121], the types of AI service-related data carried by communication connections can be naturally categorized into two major groups: AI signaling and AI data. Consequently, connection for AI embodies providing connectivity support for AI signaling and data.
@@
AI Data: 1) The AI data is used to transfer AI input data, generally referred to as input vectors for invoking an AI model or represented as inputs in an AI algorithm. 2) It is also used for transmitting AI outputs, generally called the output of the invoked AI model/algorithm. 3) It can be used for transmitting AI models, e.g., transmitting a neural network model usually means transmitting its parameter weights (which can be represented as vectors) and its network structure (which can be represented as scalars). 4) It can be further used for transmitting AI training data, which is usually huge and dedicated to the training phase of intelligent model.
For AI business scenarios and related connectivity performance requirements, 3GPP TS 22.261 (based on the conclusions of TR 22.874) also has some KPIs requirement level analysis and results for 5G systems, mainly on the impact of conventional connectivity performance metrics, such as throughput, latency, reliability, etc., in other words, the current 3GPP standard only analyzes the impact of the 5G system (mainly the CN portion) in terms of how it supports different AI services/applications from the functional and performance perspectives. From an overall end-to-end process analysis perspective, AI signaling and AI data involved in AI services/applications necessitate transmission via the air interface. However, in the current standards and architectures, no relevant modules guide the transmission process or standardize the transmission content. As a result, AI services, inherently demand that the network provide better connection support for AI based on connection.
@@ 3.1.2 Data for AI
Future mobile communication networks will continuously generate massive amounts of data with diverse types, and data is precisely one of the core production factors for AI, providing a solid data foundation for 6G’s endogenous intelligent AI services [130]. In the field of communications, several publicly available datasets have been widely utilized, such as DeepMIMO [131] and RadioML [132]. The DeepMIMO dataset, based on ray-tracing simulations of real wireless communication environments, provides MIMO channel data and is extensively applied in scenarios like channel modeling, beamforming, and user location in 6G networks. Meanwhile, the RadioML dataset encompasses radio signal data with multiple modulation schemes and is mainly used for tasks including signal modulation recognition, interference detection, and spectrum management. Additionally, the “AI/ML in 5G Challenge” dataset [133] released by ITU covers a variety of communication scenarios in 5G networks, offering crucial support for research on network optimization, traffic prediction, and resource allocation. These datasets not only lay a solid foundation for AI research in the communications field but also provide a unified benchmark for validating and comparing algorithm performances.
@@
Data for AI aims to efficiently support end-to-end data collection, transmission, storage, and sharing, addressing how to facilitate, expedite, and securely provide data to AI functions within or external to the 6G network. This endeavor strives to reduce the difficulty in data acquisition, enhance data service efficiency, and facilitate the realization of 6G AI while adhering to privacy and security laws and regulations. Depending on the potential functional scope, the data support provided by data for AI should encompass five components:
@@
Data Collection/Distribution: This component provides a basic publish-subscribe mechanism for data producers and consumers, enhancing data collection/distribution efficiency and supplying samples for large-scale training required by AI models to achieve optimal results.
@@
Data Security and Privacy: This component leverages security and privacy protection technologies to provide high-quality, trusted data services tailored to user and network needs, ensuring user and network privacy protection and data security, immutability, and traceability. By safeguarding data privacy, it also protects the security of AI models during training and data transmission.
@@
Data Analysis: This component utilizes AI models, algorithms, and knowledge to provide statistical information, predictive insights, network anomaly analysis, and optimization suggestions, enhancing the data consumption experience for internal and external network functions and related intelligent applications.
@@
Data Preprocessing: This component applies generic tools for format conversion, noise reduction, feature extraction, and other preprocessing tasks to collected data, fulfilling AI model training requirements for input data and facilitating the management of multi-dimensional complex data.
@@
Data Storage: This component stores and retrieves collected data, providing storage support for data services related to data security and privacy, data analysis, or data preprocessing, and further establishing a data foundation to support AI services.
Through standardized data service functions, a complete data service process can be established within the 6G network, providing a robust foundation for AI at the data level. Meanwhile, in view of the fact that researchers often need to construct dedicated datasets for specific application scenarios, Data for AI provides valuable guidance. Taking the construction of datasets in the field of communications as an example, we offer the following practical experiences. Firstly, data sources can be selected from collected data in real communication networks or generated simulation data using high-fidelity simulation tools (such as NS-3 or OMNeT++). Real data can reflect actual scenarios but may involve privacy and security issues; whereas simulation data offers higher flexibility and can be customized to specific scenarios according to particular requirements. Secondly, data preprocessing and labeling are essential steps. Communication data usually contains a large amount of noise, thus requiring preprocessing operations such as cleaning and normalization, and accurate labeling based on research objectives (such as channel state prediction or interference detection). Moreover, to enhance the generalization ability of the model, diverse communication scenarios and multiple network conditions should be covered as much as possible when constructing the dataset. Meanwhile, when using real user data, strict compliance with data privacy protection regulations is mandatory, and techniques such as anonymization or differential privacy should be adopted to ensure data security. Finally, we encourage researchers to make datasets publicly available under the premise of protecting privacy, in order to promote collaborative research and algorithm comparison within the field, and it is recommended to use a unified storage format and standard to facilitate data sharing and the reproducibility of research results.
@@ 3.1.3 Computation for AI
Computation is another core production factors of AI. In future network architectures, computational power will be ubiquitous across various infrastructures, including central clouds, edge clouds, network devices, and even end-user devices, providing a solid computational foundation for the diversification of AI services across all industries in 6G [134]. This computational power, along with the AI algorithms or functionalities attached, serves the network or devices to improve performance and optimize network operations but is also potentially exploited through unified interfaces to serve upper-layer applications. Consequently, computation for AI focuses on how the widely available computational power in the network can support AI services. This computational power can be broadly categorized into three types: dedicated computing power for network elements, distributed external computing power, and endogenous computing power within the network.
Network Element Dedicated Computing Power (for AI): This type of computing power typically consists of computing and storage units composed of specialized processors and programmable devices. It serves network elements in mobile communication networks (BS or CNs). It is primarily used for customized AI application services (i.e., AI4NET) to enhance communication performance or optimize network operations, like CSI feedback, channel estimation, and beamforming. Due to resource restrictions, it cannot support AI services and applications requiring large-scale computations and training. It is challenging to implement for third-party applications.
Distributed External Computing Power (for AI): This type of computing power typically exists in the form of distributed edge computing/MEC, primarily utilizing general-purpose central processing units (CPU) and potentially including high-performance graphics processing units (GPU) and programmable acceleration cards. External computing power shifts computations from centralized data centers to the edge of the access network, enabling network optimization and supporting industry applications with high computational demands and stringent latency requirements, such as video acceleration and augmented reality (AR)/VR scenarios. Due to its external nature, it requires management through unified network functions, introducing some latency that may impact extremely latency-sensitive AI services and applications. Moreover, this external overlay approach may not optimize resource utilization, hindering efficient construction, deployment, and usage of AI services.
@@ 3.1.4 Security for AI
Some security vulnerabilities remained after the design of 2G-5G systems, and conventional security measures often failed to ensure safety. In recent years, trusted endogenous security has emerged as a new security approach for 6G, characterized by four main features: collaboration, intelligent proactive defense, trustworthiness, and privacy protection. The concept of trusted endogenous security also applies to ensuring the security of the AI plane.
@@
Trustworthiness: 6G will further evolve towards trustworthiness capabilities, achieving a comprehensive endogenous security system encompassing Security (i.e., confidentiality, integrity, availability), reliability, resilience, and safety [136]. This can provide a trusted ecosystem for NET4AI. Furthermore, flexible and reliable access control [137] and effective identity authentication[138] can, to some extent, prevent unauthorized access from impacting the endogenous trusted environment of NET4AI.
Privacy Protection: The deep integration of AI, big data, and 6G networks will exacerbate the challenges of data privacy protection. NET4AI must handle massive business and personal data while preventing privacy leakage. Emerging distributed ML, while achieving distributed ubiquitous intelligence and partially avoiding privacy leakage caused by AI model training, still has security risks [139]. Therefore, NET4AI will integrate distributed ML with existing privacy protection methods at multiple points to enhance data privacy protection and build an efficient and secure data ecosystem.
@@ 3.1.5 Orchestration for AI
Fault Self-Healing and Intelligent O&M: By integrating monitoring, alerting, and troubleshooting functions, the orchestrator promptly identifies and locates faults within the network, automatically triggering recovery mechanisms. This fault self-healing capability minimizes service disruptions for AI services, enhancing service availability and reliability. Additionally, the orchestration facilitates intelligent network O&M, empowering network administrators to manage network resources and AI services more efficiently, reducing O&M costs, and improving efficiency.
@@ 3.2 Key Enabling Technologies
This section describes the key enabling technologies for NET4AI. Note that each of the five capability/service planes in future 6G network architecture as presented in Section 3.1 shall contain its own enabling technologies as we compile in Figure 15. In the following, we only focus on the ones that are closely related to the needs of AI, introducing their research background, research significance, and research status.
@@ 3.2.1 Distributed Intelligence and Federated Learning
The NET4AI architecture must address the dual demands of low-latency, high-reliability communication and computing, as well as growing concerns about privacy. Distributed intelligence and FL are the key technologies that can solve these challenges. conventional centralized architectures in wireless networks can no longer meet low-latency, high-reliability communication and computation demands. Additionally, conventional centralized architectures are inefficient in supporting the ubiquitous intelligence required by future wireless networks. This trend towards deploying intelligent decentralized elements has become apparent, with intelligence gradually shifting from centralized to distributed systems. Therefore, introducing distributed intelligent computing architectures is necessary to fully utilize the multi-dimensional data and computing resources held by user terminals and nodes, enabling intelligent connectivity.
@@
Wireless distributed intelligence refers to organizing AI tasks in a distributed manner within wireless networks and incorporating collaborative AI and ML without uploading all raw data to a central cloud. This approach alleviates the network bandwidth pressure and reduces the central cloud’s computational burden. Moreover, as the computing power of smartphones and IoT devices increases and users become more concerned about data privacy, mobile devices can handle more AI computations and training. This proximity of AI to users and data allows for faster iteration and upgrade cycles, efficiently supporting the pervasive intrinsic intelligence of 6G networks.
@@
Distributed intelligence is a key technology for achieving the intrinsic intelligence of 6G networks. It adopts a decentralized architecture, dispersing network functions and resources across multiple nodes or devices, eliminating the dependence on a single central node to control the entire network. Nodes in a distributed network possess a certain degree of autonomy, enabling them to make decisions and execute tasks independently without frequent communication or coordination with other nodes. This autonomy enhances network flexibility, allowing it to better adapt to various environmental and demand changes. Data is typically stored across multiple nodes in a distributed network, and computational tasks can be distributed among these nodes for processing. This distributed data storage and processing method improves network scalability, accelerates the training and inference processes of large-scale models, and enhances the efficiency of AI systems. Distributed intelligence also supports online learning and continuous optimization by collecting and processing data in real-time, continuously updating models, and improving AI system performance and accuracy. By leveraging the potential of large-scale distributed edge devices, distributed AI technology can overcome the bottlenecks of centralized AI, reduce communication costs, and address data privacy concerns. Distributed AI will build a new intelligent ecosystem for 6G. As shown in Figure 16, the wireless distributed intelligence architectures widely recognized in academia and industry include:
Figure 16: Federated learning, Multi-Agent RL and Split Learning
Federated Learning: FL is a typical paradigm of distributed intelligence, which enhances privacy by training AI models on local devices and only sharing model parameters instead of raw data. During training, each node trains the model locally and uploads the model weights regularly. The central node integrates these weights to form the weights of the federated model and then feeds the weights back to each node for the next round of training or local inference. FL ensures that user data remains local and only transmits weights/gradients to achieve global training, thus enhancing privacy. A typical example of implementing model training using FL in wireless networks is shown in part a of Figure 16. The entire FL process involves downloading the global model $\omega^{t}$ of the current round $t$ to local devices; Device $k$ trains model based on its own private dataset $D_{k}$ to obtain the local gradient update $\mathbf{g}^{t}_{k}$ and then uploads it; The parameter server executes the federated averaging (FedAvg) algorithm to obtain the gradient update aggregation $G^{t}$ and updates the new global model $\omega^{t+1}$ for round $t+1$. This process will be iterated until the end. In addition, FL can be further divided into two types according to the correspondence between the feature dimensions of the dataset used and the samples: horizontal FL and vertical FL.
@@
(1) Horizontal FL: Also known as sample-based FL, this can be applied in scenarios where the datasets of the various participants in FL have the same feature space but different sample spaces. Through horizontal FL, different participants can jointly train a more powerful AI model based on local data sets while protecting privacy and improving the accuracy and generalization of the model. Additionally, horizontal FL enables real-time learning on distributed edge devices, allowing for timely model parameter updates to constantly changing environments and data. This is crucial for AI applications that require quick responsiveness and strong adaptability. It also supports collaboration across devices with varying computational capabilities and data distributions, catering to the diverse needs of 6G network terminals. Personalized model training on local devices enhances user experience by offering more individualized services.
The authors of [141] took into account the actual situation of limited resources in wireless networks, and deeply analyzed the convergence of FedAvg in horizontal FL under the two assumptions of full and partial device participation, providing important theoretical support for the applicaiton of horizontal FL in wireless networks. In [142, 143], the authors furthermore represented the partial device participation as a probability distribution and eliminate the device selection bias by modifying the aggregation rule in FedAvg to ensure the unbiasedness of gradient updates. This allowed them to provide a more stringent convergence bound and is committed to the efficient implementation of horizontal FL. Based on these theoretical supports, the authors of [144] provided a variety of solutions to guide the efficient application of FL in wireless networks, and was committed to reducing the severe energy consumption problem of FL from aspects such as layering, aggregation frequency optimization, device selection, and communication compression, so as to achieve low-carbon green economic goal. In [145], the authors focued on the shortcoming of horizontal FL in dealing with the heterogeneity of device data, and introduced a personalized federation learning (PFL) method with theoretical performance derivation, providing a theoretical basis for the applicaiton of FL in multiple fields to realize personalized models. The authors of [146] categorized PFL techniques based on key challenges and strategies, highlighting key ideas and future research directions in PFL architecture, realistic benchemarks, and trustworthy methods. In [147], the authors considered the applicaiton of horizontal FL in the medical field to solve the problem of collaborative research and patient health testing among multiple hospitals, and proposed a blockchain-based orchestration ML system framework to explore the protection of medical data privacy. The authors of [148] focued on the field of IoT healthcare and proposed an intelligent medical system agent architecture that integrated horizontal FL and blockchain technology to achieve efficient and secure processing of medical data to optimize the medical diagnosis process. The authors of [149] proposed a new FL framework for cross-domain prediction for intelligent manufacturing, designed to quickly adapt to new products, processes, and applications with limited training data.
Horizontal FL has important application prospects in providing efficient and personalized AI services in wireless networks, coordinating research with multiple hospitals in healthcare, and realizing cross-application in intelligent manufacturing. However, it also has defects such as slow convergence speed and susceptibility to data heterogeneity.
@@
(2) Vertical FL: Vertical FL refers to a setup where datasets have different features but share the same samples/users. It is suitable for scenarios where there is a high overlap of users among participants but low or no overlap in feature space by extracting parts where participants have the same users but not the same features for joint training. In the training process, it is first necessary to sample align each participant’s data to obtain overlapping user data, then train the model locally and upload the parameters to the server [150]. The difference between vertical and horizontal FL is that horizontal FL emphasizes cooperation and sharing between different data sources, while vertical FL focuses on collaboration and sharing within varying levels of the same data source.
@@
The authors of [151] proposed a vertical federated recommendation system method based on clustering and latent factor model. Facing the diversity of a large number of different users, the users are clustered, and the prediction rating model and clustered user latent factor vector are reconstructed to more fully explore the relationship between users and items and improve the accuracy of user recommendations. In [152], the authors described the idea of intent-hidden vertical FL, taking the example of multi-party collaborative training of disease prediction models in medical diagnosis, and buildt a security protocol using homomorphic encryption to meet the exact needs of data privacy protection. The authors of [153] proposed a vertical federated DL based on the spatial-temporal long and short-term networks algorithm for intelligent city traffic in MEC environment. Using different sample features (i.e. different regions) in the same sample space (i.e. the same traffic scene and time range) to build a federated model to predict traffic flow, effectively solving the problems of data privacy and spatio-temporal information capture in traffic prediction, providing more accurate prediction results, and helping to optimize traffic management and improve travel efficiency.
Vertical FL can fully explore the connections between different features of data samples and is of great significance in recommendation systems and traffic prediction. However, it is still in its early stages of research and related theoretical research has not yet been perfected, so it cannot provide sound theoretical support.